# Privacy-Aware In-Context Learning for Large Language Models

### v1

@@ Discussion
Traditional differentially private generation methods, such as those employed by Amin et al. (2024), rely on metric-based mechanisms like the Sparse Vector Technique (SVT), which requires threshold computations under additive noise. These approaches typically employ distributional distance metrics such as $\ell_{1}$ to score or select candidate tokens. However, these metrics are computed over normalized probability distributions (i.e., post-softmax), where even semantically similar tokens with slightly different probabilities or indices can yield high $\ell_{1}$ distance values. This issues is especially pronounced in high-dimensional output spaces of language models, where such metrics treat tokens as orthogonal dimensions, ignore semantic similarity, and are insensitive to token ranking that are crucial for meaningful text generation.
@@
Moreover, approach by Amin et al. (2024) continues generating additional sentences beyond termination, often yielding low-quality synthetic examples. Our procedure strictly terminates when <eos> token is encountered. This guarantees that all outputs are semantically coherent and suitable for downstream tasks. Qualitative examples supporting this claim are provided in the Appendix.
@@ Appendix A Proof of Theorem 1
Lemma 3 (Advanced Composition for Exponential Mechanism  (Dwork, Roth et al. 2014)).
@@ Appendix B Hyperparameter tuning
Table 5: Values for hyperparameters explored in this work.
@@ Appendix C Empirical privacy evaluation by membership inference attacks
Table 6: Empirical privacy evaluation of our method for 1-shot ICL by MIA on Gemma 1.1 2B IT model.
While differential privacy offers theoretical privacy guarantees, empirical evaluation is also crucial (Blanco-Justicia et al. 2022). To assess the practical privacy of our DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) (Shokri et al. 2017), a practical method for measuring real-world privacy leakage.
We follow (Duan et al. 2023) to instantiate MIA in the in-context learning (ICL) setting, where the goal is to detect whether a data point was included in the LLM prompt. As expected, using true private samples leads to successful attacks. To evaluate our DP few-shot generation, we split the DBPedia dataset into member and non-member sets. Using member data, we generate 1-shot demonstrations with our DP algorithm for $\epsilon=4$ on 5 runs. For MIA, we query 50 member and 50 non-member samples, repeating for 100 trials to compute the average AUC. For the non-private baseline, we follow the same setup using actual member samples in the prompt.
Table 6 shows the MIA results. Consistent with (Duan et al. 2023), using actual private samples yields high AUC (94.20 for $\epsilon=\infty$), indicating successful attacks. In contrast, our DP approach significantly reduces the AUC, demonstrating improved membership privacy.
@@ Appendix D Effect of model size/architecture
Table 7: DBPedia classification accuracy for various LLMs. All evaluations use the same ICL setup described in §Experimental Setup.

### v3

@@ Discussion
Traditional differentially private generation methods, such as those employed by [57], rely on metric-based mechanisms like the Sparse Vector Technique (SVT), which requires threshold computations under additive noise. These approaches typically employ distributional distance metrics such as $\ell_{1}$ to score or select candidate tokens. However, these metrics are computed over normalized probability distributions (i.e., post-softmax), where even semantically similar tokens with slightly different probabilities or indices can yield high $\ell_{1}$ distance values. This issues is especially pronounced in high-dimensional output spaces of language models, where such metrics treat tokens as orthogonal dimensions, ignore semantic similarity, and are insensitive to token ranking that are crucial for meaningful text generation.
@@
Moreover, approach by [57] continues generating additional sentences beyond termination, often yielding low-quality synthetic examples. Our procedure strictly terminates when <eos> token is encountered. This guarantees that all outputs are semantically coherent and suitable for downstream tasks. Qualitative examples supporting this claim are provided in the Appendix.
@@ Appendix A Proof of Theorem 1
Lemma 3 (Advanced Composition for Exponential Mechanism  [3]).
@@ Appendix B Hyperparameter tuning
TABLE V: Values for hyperparameters explored in this work.
@@ Appendix C Empirical privacy evaluation by membership inference attacks
TABLE VI: Empirical privacy evaluation of our method for 1-shot ICL by MIA on Gemma 1.1 2B IT model.
While differential privacy offers theoretical privacy guarantees, empirical evaluation is also crucial [64]. To assess the practical privacy of our DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) [66], a practical method for measuring real-world privacy leakage.
We follow [20] to instantiate MIA in the in-context learning (ICL) setting, where the goal is to detect whether a data point was included in the LLM prompt. As expected, using true private samples leads to successful attacks. To evaluate our DP few-shot generation, we split the DBPedia dataset into member and non-member sets. Using member data, we generate 1-shot demonstrations with our DP algorithm for $\epsilon=4$ on 5 runs. For MIA, we query 50 member and 50 non-member samples, repeating for 100 trials to compute the average AUC. For the non-private baseline, we follow the same setup using actual member samples in the prompt.
Table VI shows the MIA results. Consistent with [20], using actual private samples yields high AUC (94.20 for $\epsilon=\infty$), indicating successful attacks. In contrast, our DP approach significantly reduces the AUC, demonstrating improved membership privacy.
@@ Appendix D Effect of model size/architecture
TABLE VII: DBPedia classification accuracy for various LLMs. All evaluations use the same ICL setup described in §Experimental Setup.